A car travels 450km in 4 hours. If that journey included a 20 minute rest break, what is the average speed of the car when it was actually moving? (Round answer to the nearest whole number)


1
Expert's answer
2020-09-24T11:08:56-0400

The average speed of the car when it was actually moving excludes the break time:


"s=\frac{D}{t_m-t_b}=\frac{450}{4-20/60}=123\text{ km/h}."